A workman needs to know the temperature inside the oven. He removes 2.0-lb. iron bar from the oven and places in a 1-lb aluminum container partially filled with 2.0 lb. of water. If the temperature of the water rises from 70 to 120oF, what was the oven temperature?


1
Expert's answer
2021-02-11T08:54:20-0500

Given

Specific heat of iron: 0.11 Btu/(lb·°F)

Specific heat of aluminum: 0.258 Btu/(lb·°F)


Solution

Heat from iron as it cools:


"Q_1=c_im_i(t-t_i)."

Heat to the container and water as they heat up:


"Q_2=c_wm_w(t-t_w)+c_am_a(t-t_w)."

The heat flows result in 0:


"Q_1+Q_2=0,\\
c_im_i(t-t_i)+c_wm_w(t-t_w)+c_am_a(t-t_w)=0,\\\space\\
t_i=\frac{(t-t_w)(c_wm_w+c_am_a)+c_im_it}{c_im_i}=633°\text{F}."